OpenStudy (anonymous):

can someone google this answer and get the yahoo answer Landon is standing in a hole that is 5.1 ft deep. He throws a rock, and it goes up into the air, out of the hole, and then lands on the ground above. The path of the rock can be modeled by the equation y = -0.005x2 + 0.41x - 5.1, where x is the horizontal distance of the rock, in feet, from Landon and y is the height, in feet, of the rock above the ground. How far horizontally from Landon will the rock land?

OpenStudy (anonymous):

shabeer y u blok me

OpenStudy (anonymous):

You just keep on chatting

OpenStudy (anonymous):

@abdul shabeer im sorry can u unblok me

OpenStudy (anonymous):

@abdul shabeer click my name unblock user

OpenStudy (amistre64):

can someone google ....... that sounds like an odd request

OpenStudy (anonymous):

do u know how

OpenStudy (amistre64):

how to google? yes

OpenStudy (anonymous):

how to answer the question

OpenStudy (amistre64):

yeah, since the rock follows the path given then you just need to determine what teh x intercepts are and weed out any spurious results

OpenStudy (anonymous):

so can u solve it

OpenStudy (amistre64):

i can solve it yes.

OpenStudy (anonymous):

can u help me

OpenStudy (amistre64):

i can, but i need to know where it is that you are getting lost in this.

OpenStudy (anonymous):

what do i need to do to get the answer

OpenStudy (amistre64):

you need to know the quadratic formula; that would be the simplest way to approach this

OpenStudy (anonymous):

ax^2 + bx + c = 0

OpenStudy (amistre64):

\[y = ax^2 + bx + c\] y = -0.005x2 + 0.41x - 5.1\[\begin{matrix}a=-0.005&b=0.41&c=-5.1\end{matrix}\] the quadrtic formula is set up so that:\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (amistre64):

since we know all the values, we simply plg them into their right spots and calculate: \[x=\frac{-0.41\pm\sqrt{0.41^2-4(-0.005)(-5.1)}}{2(-0.005)}\]

OpenStudy (amistre64):

one of these values will most likely be useless to the situation; so we need to exclude it in the results as extraneous

OpenStudy (anonymous):

33.35 ft

OpenStudy (amistre64):

lol, ill trust you on that :)

OpenStudy (anonymous):

is it right

OpenStudy (amistre64):

dunno, id use a calculator to come up with it if i had one

OpenStudy (anonymous):

i dnt

OpenStudy (anonymous):

so...

OpenStudy (anonymous):

do u know what its round

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=-0.005x%5E2%2B0.41x-5.1%3D0 66.71 would be my assumption

OpenStudy (anonymous):

thanks that all u had to say in the beg.

OpenStudy (amistre64):

no, it wasnt. http://openstudy.com/code-of-conduct "Don’t post only answers - guide the asker to a solution."

OpenStudy (anonymous):

idc

OpenStudy (amistre64):

you should, otherwise all your doing is abusing this site to get free answers ...

OpenStudy (anonymous):

ok and.....

OpenStudy (anonymous):

if i was...

OpenStudy (amistre64):

that should be reason enough in and of itself. but as long as you are not harassing people to get an answer; then its all good :)

OpenStudy (anonymous):

dunno, id use a calculator to come up with it if i had one u knew it just like just so happen u assumed right! right....

OpenStudy (anonymous):

im not harassing

OpenStudy (amistre64):

and im not suspending ....

OpenStudy (anonymous):

not suspending what..................................